in an experiment the voltmeter reads 12v the ammeter 4 amp a mettalic block of mass 1kg os heated in 300s which cause a raise in temprature by 16 degree Celsius calculate the specific heat capacity of the metal
1
Expert's answer
2018-10-08T15:00:09-0400
The energy balance gives
work done=heat
IVτ=cmΔt
Thus
c=IVτ/mΔt
=(4×12×300)/(1×16)=900 J/(kg∙℃)